OpenStudy (anonymous):

I NEED HELP WITH 5 MATH PROBLEMS: step by step!!! The left has to equal the right. 1.) cot(x)sec^4x = cot(x) + 2 tan(x) + tan^3(x)

OpenStudy (anonymous):

cot(x)sec^4x=cotx ((sec^2x)^2)=cotx(1+tan^2x)^2 =cotx(1+2(tanx)^2+(tanx)^4)=1/tanx (1+2(tanx)^2+(tanx)^4) =1/tanx +2 tanx + (tanx)^3=cot(x) + 2 tan(x) + tan^3(x)

OpenStudy (anonymous):

Okay Thnx> Next problem is: (sinx)(tanx cosx - cotx cosx) = 1-2 cos^2x